# Electricity Price-Aware Scheduling of Data Center Cooling

### v1

@@ Abstract
Data centers are becoming a major consumer of electricity on the grid, with cooling accounting for about 40% of that energy. As electricity prices vary throughout the day and year, there is a need for cooling strategies that adapt to these fluctuations to reduce data center cooling costs. In this paper, we present a model for electricity price-aware cooling scheduling using a Markov Decision Process(MDP) framework to reliably estimate the cooling system operational costs and facilitate investment-phase decision-making. We utilize Quantile Fourier Regression (QFR) fits to classify electricity prices into different regimes while capturing both daily and seasonal patterns. We simulate 14 years of operation using historical electricity price and outdoor temperature data, and compare our model against heuristic baselines. The results demonstrate that our approach consistently achieves lower cooling costs. This model is useful for grid operators interested in demand response programs and data center investors looking to make investment decisions.
@@ II-G Optimization Problem
We now introduce the linear program (3) to compute the optimal cooling policy for the data center (for a detailed explanation see [16]).
@@ III-C Policy Visualization
Figure 2: Optimal daily cooling action plan under an electricity price model with four independent chillers. Colors represent the number of chillers in use: red—all four chillers active, orange—two chillers active, yellow—no chillers active. Upper panel: Four-regime. Lower panel: Eight-regime.
@@ IV Numerical Results
In this section we validate the performance of our proposed QFR-based MDP model through simulations, using real-time electricity price data from 2011 to 2024 from ISO New England, and outside temperature data from the Boston, MA, station. The simulation is conducted in-sample from 2011 to 2020 and out-of-sample from 2021 to 2024. In order to do that, we introduce two well-known heuristic methods of cooling as baselines.
@@
Fig. 3 shows the indoor temperature on a sample date (July 15th, 2024) under different policies. In all cases, the system maintains temperature within safe bounds, 18–27°C. The Greedy method holds the temperature at the upper limit $T_{\max}=$27°C. The Fixed rule aggressively cools in the middle of the night (2 AM - 3 AM) and turns off during peak hours (3 PM - 7 PM). In contrast, our proposed QFR-MDP method dynamically adapts cooling to stay within the comfort range based on real-time prices. Fig. 4 compares the energy cost of cooling the data center under different policies in a simulation across multiple years. The QFR-MDP method consistently achieves lower energy costs compared to both baselines. This stems from the inability of the other two methods to react to stochastic variability in electricity prices. These results demonstrate the potential of QFR-MDP scheduling for electricity price-aware, thermally safe data center operations.
@@
Figure 4: Annual cooling cost comparison across control methods from 2011 to 2024. The QFR-MDP approach consistently achieves lower cooling costs compared to the Greedy and Fixed Rule baselines.
@@ V Conclusion
We present an MDP-based model for electricity price-aware cooling in a data center, utilizing QFR fits to categorize real-time prices into different price regimes. These fits are time-dependent, enabling our model to capture both the diurnal and seasonal patterns in the electricity price. This model is specifically designed for investment phase decisions by providing an estimate of the operational costs and plan for cooling a data center. It offers an accurate estimate of optimized cooling costs well before the construction stage of a data center which is important for both engineering design and financial planning.
Through 14 years of simulations using real-time temperature and electricity price data, we demonstrate that our model consistently has lower cooling costs compared to other heuristic baselines. Future work may explore the co-optimization of price-aware cooling and workload scheduling in data centers, as well as the integration of energy storage systems such as batteries with the QFR-MDP model.

### v2

@@ Abstract
Data centers are becoming a major consumer of electricity on the grid, with cooling accounting for about 40% of that energy. As electricity prices vary throughout the day and year, there is a need for cooling strategies that adapt to these fluctuations to reduce data center cooling costs. In this paper, we present a model for electricity price-aware cooling scheduling using a Markov Decision Process (MDP) framework to reliably estimate the cooling system’s operational costs and facilitate investment-phase decision-making. We utilize Quantile Fourier Regression (QFR) fits to classify electricity prices into different regimes while capturing both daily and seasonal patterns. We simulate 14 years of operation using historical real-time electricity price and outdoor temperature data, and compare our model against heuristic baselines. The results demonstrate that our approach consistently achieves lower cooling costs. This model is useful for grid operators interested in demand response programs and data center investors looking to make investment decisions.
@@
$c_{t,s,a}$
Immediate cost of being in state $s$ at time $t$ and taking action $a$ (\$).
$P_{t,s,s^{\prime},a}$
Transition probability from $s=(\theta,p)$ to $s^{\prime}=(\theta^{\prime},p^{\prime})$ under action $a$ at time $t$.
@@ II-G Optimization Problem
We now introduce the linear program (3) to compute the optimal cooling policy for the data center.
@@ III-C Policy Visualization
Figure 2: Optimal daily cooling action plan under the QFR-MDP electricity-price model for a data center with four independent chillers (example day: July 15). Colors indicate the number of active chillers: red—all four chillers active, orange—two chillers active, yellow—no chillers active. Upper panel: Four-regime. Lower panel: Eight-regime.
@@ IV Numerical Results
In this section, we validate the performance of our proposed QFR-based MDP model through simulations, using real-time electricity price data from 2011 to 2024 from ISO New England, and outside temperature data from the Boston, MA, station. The simulation is conducted in-sample from 2011 to 2020 and out-of-sample from 2021 to 2024. In order to do that, we introduce two well-known heuristic methods of cooling as baselines.
@@
Fig. 3 shows the indoor temperature on a sample date (July 15th, 2024) under different policies. In all cases, the system maintains temperature within safe bounds, 18–27°C. The Greedy method holds the temperature at the upper limit $T_{\max}=$27°C. The Fixed rule aggressively cools in the middle of the night (2 AM - 3 AM) and turns off during peak hours (3 PM - 7 PM). In contrast, our proposed QFR-MDP method dynamically adapts cooling to stay within the comfort range based on real-time prices. Fig. 4 compares the energy cost of cooling the data center under different policies in a simulation across multiple years of summers. The QFR-MDP method consistently achieves lower energy costs compared to both baselines. This stems from the inability of the other two methods to react to stochastic variability in electricity prices. These results demonstrate the potential of QFR-MDP scheduling for electricity price-aware, thermally safe data center operations. The fact that the eight-state system achieves lower costs than the four-state system in most summers suggests that using finer bins can lead to more effective policies in terms of average cost.
@@
Figure 4: Relative improvement in cooling cost across control methods from 2011 to 2024 summers. The QFR-MDP approach consistently achieves lower cooling costs compared to the Greedy as a baseline.
@@ V Conclusion
We present an MDP-based model for electricity price-aware cooling in a data center, utilizing QFR fits to categorize real-time prices into different price regimes. These fits are time-dependent, enabling our model to capture both the diurnal and seasonal patterns in the electricity price. This model is specifically designed for investment phase decisions by providing an estimate of the operational costs and plan for cooling a data center. It offers an accurate estimate of optimized cooling costs well before the construction stage of a data center, which is important for both engineering design and financial planning.
Over 14 years of summer simulations using real-time electricity price data, we show that our model consistently has lower cooling costs compared to other heuristic baselines. Future work may explore the co-optimization of price-aware cooling and workload scheduling. Another direction is to extend our framework to incorporate broader sustainability objectives, such as reducing carbon emissions from electricity use while considering fairness (see [16, 17]).